OpenStudy (anonymous):

Solve cos2x=1/√(2) Please show steps!

OpenStudy (raden):

what are inteval for x?

OpenStudy (anonymous):

\[0\le x \le2\pi\]

OpenStudy (across):

Well, you know that\[\cos\left(\pi/4\right)=\frac{\sqrt2}{2}=\frac{1}{\sqrt2}.\]Immediately you know that\[x=\frac{\pi}{8}+\pi k,\]where \(k\in\mathbb{Z}\) are the solutions.

OpenStudy (anonymous):

that really helped, thanks!

OpenStudy (raden):

i think the answer not complete... how about u across, if x=7pi/8 ? right to, isn't it?